OpenStudy (anonymous):

Find the vertex, focus, directrix, and focal width of the parabola. -(1/12)x^2=y

OpenStudy (anonymous):

you are stuning beuty

OpenStudy (anonymous):

i finished this one but i have another problem :)

OpenStudy (solomonzelman):

Okay, good :)

OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±6) and asymptotes at y = ±3/2.x.

OpenStudy (solomonzelman):

I am not good at finding it backwards -:(

OpenStudy (solomonzelman):

Actually...

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{x^2}{36}-\frac{y^2}{not~~sure.} =1 }\)

OpenStudy (anonymous):

that's what i was stuck on xD

OpenStudy (anonymous):

its either 16 or 4

OpenStudy (solomonzelman):

ohh... the asymptotes are at y=±3/2. So a=±2 and b=±3 So it should be \(\LARGE\color{black}{ \frac{x^2}{4}+\frac{y^2}{9} =1 }\)

OpenStudy (solomonzelman):

it is not minus between the fractions, confused with an ellipse.

OpenStudy (anonymous):

oh wow im totally wrong

OpenStudy (anonymous):

yes i under stand what i did wrong haha thank you

OpenStudy (solomonzelman):

But just looking at the asymptotes, I found it... I am not sure ... I think the vertex contradicting I am not sure about this. Sorry

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

i have another question but i should make a new question for it so medals can be given out haha

OpenStudy (solomonzelman):

I think I can do it, give me a last chance.

OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±9) and foci at (0, ±10).

OpenStudy (anonymous):

(y^2/81)-(x^2/100)=1

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

its horizontal so a^2 the higher value would be on the right

OpenStudy (solomonzelman):

vertex is at (0,±6) thus we get \(\LARGE\color{black}{ \frac{x^2}{36} }\), and so far we can say \(\LARGE\color{black}{ \frac{x^2}{36}+\frac{y^2}{...}=1 }\) The asymptotes are at, y=±3/2 x So, y=±3/2 x → y=±6/4 x \(\LARGE\color{black}{ \frac{x^2}{36}+\frac{y^2}{16}=1 }\)

OpenStudy (solomonzelman):

I think this is it.

OpenStudy (solomonzelman):

Switch 36 and 16.

OpenStudy (anonymous):

ohhhh b^2=a^2-c^2

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{x^2}{16}+\frac{y^2}{36}=1 }\) because 4 is a, and 6 is b.

OpenStudy (anonymous):

so b would be 19

OpenStudy (anonymous):

y^2/81 - x^2/19 = 1

OpenStudy (anonymous):

you're answer is not achoice

OpenStudy (solomonzelman):

for hyperbola, dang.. it is \(\LARGE\color{black}{ \frac{x^2}{16}-\frac{y^2}{36}=1 }\) for the second question. For the last one, yes, I think you are correct.

OpenStudy (solomonzelman):

I am very confusing often times.

OpenStudy (anonymous):

ahahahha yessss thats the one for last one